# Multimodal Banking Dataset: Understanding Client Needs through Event Sequences

### v1

@@ 3.2.1 Campaigning task
In this task, it is required to predict the customer’s propensity to purchase four different products in the next month (Fig. 1) given sequences of transactions, geo locations, and dialogues from the beginning of this month. Solutions to this problem are used to plan marketing campaigns and prepare sales communications through various communication channels with the client.
We implement an out-of-fold validation protocol to conduct our experiments. The entire client dataset is divided into five folds: four folds are used for training, while the remaining fold is designated for testing. Both the training and testing sets are made publicly available alongside our dataset to enable future researchers to compare performance metrics.
@@
The baseline methods outlined earlier are applied as follows. First, we train our models using the training set. Considering the temporal structure of our target, we compute the embedding of each client’s history up to one month, focusing on the presence of the target product (Fig. 1). We then evaluate the model using the multi-label classification metric ROC AUC across the 12 months of 2022 and for four binary product labels.
3.2.2 Multimodal Matching
@@
Multimodal matching [41] involves aligning and comparing different modalities to identify meaningful relationships or connections between them. Frequently, data from multiple sources for the same client are matched using predefined rules or heuristics, which may not always yield optimal results. To enhance the accuracy of this process, specialized identification algorithms are required to more precisely compare different modalities.
@@
For the matching task, we employed a framework analogous to CLIP [31]. We utilized GRU encoders to embed pairs of samples from two input modalities, labeling them as either positive matches (i.e., data from the same client) or negative matches (i.e., data from different clients). The model was trained using the InfoNCE loss function [13], which maximizes similarity for positive pairs while minimizing it for negative pairs. To assess the model’s performance, we used Recall@1, Recall@50 and Recall@100 metrics.
4 Experiments
@@
Our models, experiments, and training procedures were implemented in Python, leveraging PyTorch and PyTorch Lightning for deep learning tasks, and PySpark for distributed data processing. We trained the neural networks using NVIDIA V100 GPU, while the boosting models were trained on computational clusters equipped with 600 cores. The reported experiments, including extensive hyperparameters optimization, required approximately 500 hours of computation time.
4.1 Model and training hyperparameters
@@
We employed the unsupervised baseline methods (CoLES, TabGPT, TabBERT, Aggregation) with default hyperparameters from the pytorch-lifestream framework. The PyTorch implementation of the Adam optimizer was utilized, with an initial learning rate of 0.001, coupled with the StepLR scheduler. Models were saved based on the lowest validation loss or the highest validation unsupervised metrics [33], evaluated after each training epoch, with training of 15 epochs.
We employed 24-dimensional embeddings for categorical features and clipped the number of categories for features with many unique values. We applied either an identical mapping or a logarithmic transformation for numerical features. We used the gradient boosting algorithm available in PySpark ML for downstream tasks.
4.2 Campaigning task
In this Subsection, we provide experimental results for the campaigning task in unimodal and multimodal baselines.
@@ 4.2 Campaigning task
One of the main objectives of our paper is to provide a real data benchmark to facilitate the development of multimodal algorithms. To achieve this, we must demonstrate that an algorithm outperforming another on our public benchmark will similarly outperform it on real data.
@@
Table 1 shows that the transaction modality plays the most crucial role in achieving accurate classification. In contrast, dialogues and geostream, when used in a unimodal setting, performed only slightly better than a random estimator. However, as shown in Table 2, the predictive performance improved significantly in the multimodal setting through the integration of additional modalities. The overall trend indicates a consistent improvement in validation metrics as more modalities are incorporated. Specifically, the multimodal late fusion approach enhanced predictive accuracy by 1–2% when adding a single data source to the transaction stream, and by up to 3% when utilizing all available data sources.
Models trained on public datasets generally exhibit higher performance scores, which is an anticipated outcome of the data anonymization process. Nevertheless, the relative ranking of different methods remains consistent across both MBD and private datasets. Fig. 7 in Appendix B demonstrates a strong correlation in performance metrics between the public and private datasets, with a Kendall-tau correlation coefficient of 0.94.
@@
Comprehensive results for each modality and all possible fusion combinations of multiple modalities are provided in Appendix B. We denote our modalities: geostream, transactions, and dialogues are represented as Geo, Trx, and Dialogs, respectively. We specify the method with the modality to indicate that a particular method was applied. For example, if embeddings from the CoLES model are used on transactions, we denote this as TrxCoLES.
| Modalities | Methods | MBD | Private dataset |
| Dialog | Mean | 0.595 0.002 | 0.604 0.001 |
| Last | 0.586 0.001 | 0.590 0.001 | |
| Supervised | 0.540 0.012 | 0.538 0.010 | |
| Geo | Aggregation | 0.555 0.001 | 0.554 0.001 |
| TabGPT | 0.621 0.003 | 0.622 0.001 | |
| TabBERT | 0.603 0.002 | 0.596 0.002 | |
| CoLES | 0.598 0.004 | 0.601 0.004 | |
| Supervised | 0.566 0.002 | 0.563 0.002 | |
| Trx | Aggregation | 0.783 0.002 | 0.791 0.017 |
| TabGPT | 0.802 0.002 | 0.796 0.000 | |
| TabBERT | 0.762 0.004 | 0.754 0.011 | |
| CoLES | 0.773 0.002 | 0.772 0.002 | |
| Supervised | 0.819 0.002 | 0.813 0.002 |
| Methods | Modalities | MBD | Private dataset |
| CoLES | Trx | 0.773 0.002 | 0.772 0.002 |
| Trx + Geo | 0.775 0.002 | 0.772 0.002 | |
| Trx + Dialog | 0.781 0.001 | 0.776 0.002 | |
| Trx + Dialog + Geo | 0.783 0.001 | 0.777 0.002 | |
| TabGPT | Trx | 0.802 0.001 | 0.796 0.000 |
| Trx + Geo | 0.800 0.001 | 0.796 0.001 | |
| Trx + Dialog | 0.810 0.002 | 0.805 0.001 | |
| Trx + Dialog + Geo | 0.808 0.001 | 0.802 0.001 | |
| Supervised | Trx | 0.819 0.002 | 0.813 0.002 |
| Trx + Geo | 0.819 0.001 | 0.815 0.001 | |
| Trx + Dialog | 0.821 0.0006 | 0.817 0.0006 | |
| Trx + Dialog + Geo | 0.824 0.001 | 0.821 0.001 | |
| TabBERT | Trx | 0.762 0.004 | 0.754 0.011 |
| Trx + Geo | 0.764 0.004 | 0.756 0.011 | |
| Trx + Dialog | 0.773 0.003 | 0.765 0.009 | |
| Trx + Dialog + Geo | 0.775 0.003 | 0.766 0.009 |
4.3 Multimodal matching task
In this subsection, we present the results of our proposed multimodal matching benchmark, summarized in Table 3, which includes both transactions and dialogues. Detailed results for other modalities can be found in Appendix B, in Table 9. In this evaluation, we report recall@1, recall@50, and recall@100 for each modality, measured in both directions. For example, in the case of the transaction and geostream pair, we compute both Trx2Geo and Geo2Trx, allowing for an evaluation of alignment from both perspectives.
@@ 4.3 Multimodal matching task
Our analysis reveals considerable variation in performance across different modality pairs. Specifically, dialogue data consistently exhibits weaker matching performance compared to other modalities, such as transactions and geostream, which demonstrate significantly stronger alignment. This disparity suggests potential limitations within the dialogue modality, indicating that it may offer less complementary or aligned information. Alternatively, the unique structure of dialogue data may necessitate more sophisticated or specialized techniques for effective integration with other modalities.
These findings provide important insights for the advancement of multimodal matching methods, underscoring the need for further refinement.
| Recall@1 | Recall@50 | Recall@100 | |
| Trx2Geo | 0.006 0.0003 | 0.1967 0.002 | 0.303 0.004 |
| Geo2Trx | 0.004 0.0003 | 0.1624 0.002 | 0.262 0.004 |
5 Related Work
5.1 Financial data
The multitude of services and processes in banks generates a variety of data that can be considered as modalities. Early works such as [29, 26] use feature processing techniques that remove multimodal complexity and present data in tabular form. The Amex dataset[3] improves information content and complexity. Here, a wide range of different financial aggregates are presented as a sequence of historical slices. The development of deep learning methods has led to the ability to work with complexly structured data, such as sequences of events. Quite a few datasets [1, 5, 2], mostly monomodal, presented mainly at ML competitions. To work with such data, both supervised[7, 9, 36] and unsupervised methods[30, 8, 32] are used. Multimodal financial sequential data proposed in DataFusion2022 [4]. There are two sequential modalities, transaction and web clickstream, and two downstream tasks: matching and education level prediction. However, this is an extremely small dataset of 17K clients, and no accurate baseline model is available.
5.2 Other event sequence domains
Temporal point process[28, 40] model streams of discrete events in continuous time by constructing a neurally multivariate point process. The authors use a large collection of datasets from different types of modalities: Media (Retweets[39], MemeTrack[24], Amazon[6], IPTV[25]), Medical (MIMIC-II[22]), Social(Stack Overflow[24], Linkedin[38]) and Financial (Transaction[15]) data. All datasets are independent, and each one is single-modal. EventStreamGPT [27] uses multimodal medical record datasets, MIMIC-IV[21]. The authors propose a GPT-like approach for continuous-time event sequences. The structure of this dataset is close to the MDB. But it must be emphasized that financial data, unlike medical data, contains longer chains of events, more regular patterns, and individual transactions are less informative. These features may influence the quality of the methods used.
5.3 Geostream and dialogues
Geodata is used for various tasks. One of the uses of geo is a visualization of analytics on a map [17]. However, geo is used here not as a separate modality but as additional tags to the mainstream of tweets. Mobile marketers[12] use geo-targeting for pricing and send personalized recommendations. In [35] geo hashes are used for user mobility detection and prediction.
We encode our dialogue entries via a pretrained NLP Model. Previously, such models have already been used[18] for text anonymization tasks. Embeddings preserve the meaning of the text, which was shown in[34]. Pre-trained text embeddings can capture text sentiment and improve text-to-speech models[19].
6 Limitations
The study used data from only one company, which may limit the generalizability of the results to other organizations. In addition, the data was subject to de-identification, which limits the possibility of using models trained on this dataset outside of it. Also, the data analysis results can not be generalized. In other words, based on this dataset, it is impossible to draw conclusions regarding specific regions and market characteristics or perform deep text analytics. However, within the dataset, the data is consistent, which allows us to draw correct conclusions about the performance of multimodal or unimodal methods for working with sequences. It is also worth noting that the study was conducted on a sample of clients of a certain segment who had the opportunity to purchase certain products and does not cover all possible groups of consumers.
7 Ethics and Societal Implications
Our dataset is designed to facilitate the development of algorithms for analyzing complex event data, with the goal of enhancing the accuracy and efficiency of predictive models. Our event data does not contain characteristics such as gender, age, or race, which can be used to discriminate against certain groups of people. This makes the models trained on our dataset more ethical. However, it is important to note that companies must comply with legal grounds and obtain customer consent before collecting and analyzing additional data. It is also worth mentioning the need to ensure the security of data storage and transmission to prevent leakage of confidential information. Failure to comply with these conditions may result in privacy violations.
8 Conclusion and future work
In this paper, we introduced the first large-scale publicly available multimodal banking dataset, MBD333https://huggingface.co/datasets/ai-lab/MBD, which contains anonymized sequential data (bank transactions, geo position, and technical support dialogues) for 1M bank clients. It can be used to create industrial multimodal models for event sequences to predict client needs. Using this dataset, we present the novel benchmark based on MBD with two practically important tasks: campaigning (purchase prediction for the next month) and matching different modalities of clients. Our experimental results demonstrate that even a simple fusion of multiple data sources improves the overall quality compared to single-modal baselines (Table 2). Also, we show that anonymization does not significantly impact the algorithms’ performance, so our dataset is appropriate for choosing the best model that can be implemented in production scenarios. As a result, our work opens new horizons for developing practically relevant large-scale multimodal algorithms for event sequences.
In the future, expanding the range of tasks with our dataset is possible. For example, it is possible to estimate the multimodal techniques of the next-event prediction. Moreover, or baseline methods are based on late fusion, so it is very important to study more sophisticated models that consider possible mutual influence between modalities.
Appendix A Dataset Statistics

### v2

@@ 4.1. Datasets and downstream tasks
1. MBD. For our dataset, we introduce a campaigning downstream task (Pavlova, 2024; Balyemah et al., 2024; Hendriksen et al., 2020). In this task, it is required to predict the customer’s propensity to purchase four different products in the next month (Fig. 1) given sequences of transactions, geo locations, and dialogues from the beginning of this month. Solutions to this problem are used to plan marketing campaigns and prepare sales communications through various channels with the client.
@@
The baseline methods outlined in the following section are applied. The multi-label classification models are trained using our training set. Considering the temporal structure of our target, we compute the embedding of each client’s history for up to one month, focusing on the presence of the target product (Fig. 1). We then evaluate the model using the mean ROC AUC across four binary product labels over the 12 months of 2022.
2. Datafusion. In this dataset, the downstream task is to predict the higher education attainment of bank clients. It involves analyzing two modalities (transaction histories and clickstream data) to accurately infer client’s educational background. The task is formulated as a binary classification problem, with 75% of the labels corresponding to clients with higher education.
3. Alphabattle. We incorporate the large unimodal Alphabattle dataset alongside the multimodal datasets into our benchmark. This inclusion of data from various financial institutions aims to support more robust and reliable conclusions. In this dataset, the downstream task estimates the probability of a customer defaulting based on their historical card transaction behavior. This task is framed as a binary classification problem, with 2.7% of the labels representing clients who have defaulted. Although the dataset is unimodal, the downstream task remains highly relevant for financial institutions, offering critical insights into credit risk assessment in improving decision-making processes related to customer management and financial strategies.
@@
Multimodal matching. For the multimodal datasets, MBD and Datafusion, we propose a downstream task of multimodal matching (Zong et al., 2023; Huang et al., 2021; Taghibakhshi et al., 2023; Phan et al., 2017). It involves aligning and comparing modalities to identify meaningful relationships or connections. Data from multiple sources for the same client are frequently matched using predefined rules or heuristics, which may not always yield optimal results. To improve its accuracy, specialized identification algorithms are required to compare modalities more precisely.
@@
For the matching task, we employ a framework analogous to CLIP (Radford et al., 2021). We utilize GRU encoders to embed pairs of samples from two input modalities, labeling them as either positive (i.e., data from the same client) or negative matches (i.e., data from different clients). The model is trained using the InfoNCE loss function (Chen et al., 2020), which maximizes similarity for positive pairs while minimizing it for negative pairs. We use Recall@1, Recall@50, and Recall@100 metrics to assess the model’s performance.
4.2. Methods
To establish performance baselines, we implement several widely adopted architectures. Our approach prioritizes unsupervised and semi-supervised methods (Balestriero et al., 2023), enabling the training of a general-purpose encoder on unlabeled sequential data. Additionally, we incorporate supervised methods that allow for the immediate training of the encoder in a fully supervised manner.
4.2.1. Unimodal approaches
The following techniques are implemented in our benchmark to extract features from data. First, Aggregation Baseline that contains hand-crafted aggregation statistics (Babaev et al., 2022). Events are represented either numerically, such as transaction amount, or categorically, like event types. For numerical attributes, we apply aggregation functions (e.g., sum, mean, std) across all events in a sequence. Categorical attributes are grouped by unique values, aggregating numerical attributes using functions like count or mean.
Second, CoLES (Contrastive Learning for Event Sequences), a self-supervised contrastive model (Babaev et al., 2022) specially developed to obtain representations of such event sequences as bank transactions. The sequence encoder is a GRU (Gated Recurrent Unit) with a hidden size 256. Despite its simplicity, CoLES continues to demonstrate top-level performance in event sequence analysis (Osin et al., 2024).
Finally, two Tabular Transformers from IBM (Padhi et al., 2021): 1) TabBERT that adapts BERT to event sequences such as bank transactions, and 2) TabGPT that was initially proposed to generate synthetic tabular sequences. Both models extract 256-dimensional embeddings of an input event sequence. After that, we pool output embeddings of the client in result embedding of size 1024, calculating min, max, mean, and std.
To obtain a representation of a sequence of dialogues, we borrow conventional techniques to aggregate the sequence of embeddings of each client’s dialog: 1) mean pooling of all embeddings, and 2) use only the most recent embedding for the date of interest.
For unimodal supervised methods (Supervised RNN), we utilize GRU architectures (Babaev et al., 2019) with a hidden size of 32. The models are trained in a multi-label setting using binary cross-entropy (BCE) loss, ensuring effective optimization for tasks with multiple targets.
4.2.2. Multimodal approaches
To explore the potential of multimodal processing for event sequence analysis, we compare several fusion techniques: Blending, Late, and Early Fusion. Blending computes a weighted sum of class posterior probabilities from individual single-modal classifiers, effectively combining the predictions from each modality.
In Late Fusion, embeddings from all data sources are concatenated and fed into a classifier. This technique allows the model to learn interactions between modalities after they have been individually processed (Huang et al., 2020). In supervised Late Fusion, we utilize separate GRU encoders for each modality, concatenating their embeddings to form a unified representation (Supervised RNN).
Early Fusion combines representations from multiple modalities at the initial stages of the model, enabling joint processing of multimodal data. We employ the CrossTransformer approach (Zhang and Yan, 2023), which utilizes a cross-attention mechanism to integrate information across modalities efficiently. In our experiments, this method is applied within a supervised learning framework.
4.2.3. LLM approaches
In addition to the methods described above, we also explore the use of Large Language Models (LLMs) for multimodal data analysis. Leveraging their ability to process sequential data through universal text-based representations, LLMs offer a versatile approach to analyzing events across different modalities.
We used the X-separated serialization (Fang et al., 2024) to represent transactions in text format. X-separated serialization formats the data such that each transaction is represented as a separate line (row), with individual features separated by tabs (columns). In an unimodal setup, we limited transactions to 256 events and four features (event_time, event_type, amount, src_type3). In the multimodal setup, we used 128 transactions and 64 geo events. Serialized transactions and geo hashes are placed one after the other and separated. We used a special delimiter to separate modalities.
We applied LLMs on serialized data, considering each user as a distinct document. Transactional sequence embedding is obtained from the last hidden layer’s last token. No extra model pretraining or prompting has been applied. These embeddings were used to solve downstream tasks for each target. We also used PCA on the embeddings to reduce their dimensionality to 256, matching the baseline embeddings. Our findings indicate that applying PCA to text embeddings reduces dimensionality and improves model predictive performance on downstream tasks. We used a popular AutoML library, LightAutoML (Vakhrushev et al., 2021) to solve downstream tasks. It provides a fully automatic pipeline for model selection and hyperparameter tuning.
| Model | MBD | Datafusion | Alphabattle | ||
| Transactions | Geostream | Transactions | Clickstream | Transactions | |
| Aggregation | 0.783 0.002 | 0.595 0.002 | 0.793 0.013 | 0.537 0.018 | 0.785 0.0010 |
| CoLES | 0.773 0.002 | 0.598 0.004 | 0.784 0.012 | 0.641 0.013 | 0.793 0.0005 |
| TabBERT | 0.762 0.004 | 0.603 0.002 | 0.762 0.014 | 0.590 0.026 | 0.778 0.0003 |
| TabGPT | 0.802 0.002 | 0.621 0.003 | 0.766 0.013 | 0.618 0.016 | 0.775 0.0010 |
| Supervised RNN | 0.819 0.002 | 0.540 0.012 | 0.712 0.016 | 0.563 0.011 | 0.792 0.0030 |
| Dataset | Modalities | CoLES | TabGPT | TabBERT | Supervised RNN |
| MBD | Trx | 0.773 0.002 | 0.802 0.001 | 0.762 0.004 | 0.819 0.002 |
| Trx + Geo | 0.775 0.002 | 0.800 0.001 | 0.764 0.004 | 0.819 0.001 | |
| Trx + Dialog | 0.781 0.002 | 0.810 0.002 | 0.773 0.003 | 0.821 0.0006 | |
| Trx + Dialog + Geo | 0.783 0.002 | 0.808 0.001 | 0.775 0.003 | 0.824 0.001 | |
| Private dataset | Trx | 0.772 0.002 | 0.796 0.000 | 0.754 0.011 | 0.813 0.002 |
| Trx + Geo | 0.772 0.002 | 0.796 0.001 | 0.756 0.011 | 0.815 0.001 | |
| Trx + Dialog | 0.776 0.002 | 0.805 0.001 | 0.765 0.009 | 0.817 0.0006 | |
| Trx + Dialog + Geo | 0.777 0.002 | 0.802 0.001 | 0.766 0.009 | 0.821 0.001 | |
| Datafusion | Trx | 0.784 0.012 | 0.766 0.013 | 0.762 0.014 | 0.712 0.015 |
| Trx + Click | 0.785 0.011 | 0.766 0.011 | 0.761 0.012 | 0.703 0.008 |
5. Experiments
@@
Our models, experiments, and training procedures were implemented in Python, leveraging PyTorch and PyTorch Lightning for deep learning tasks, and PySpark for distributed data processing. We trained the models on NVIDIA V100 GPU, while the gradient boosting models were trained on computational clusters equipped with 600 CPU cores.
5.1. Model and training hyperparameters
@@
We employ the unsupervised baseline methods (CoLES, TabGPT, TabBERT, Aggregation) with default hyperparameters from the pytorch-lifestream framework. The Adam optimizer is utilized, with an initial learning rate of 0.001, coupled with the StepLR scheduler. Models are saved based on the lowest validation loss or the highest validation unsupervised metrics (Tsitsulin et al., 2023), evaluated after each training epoch, with training of 15 epochs. We employ 24-dimensional embeddings for categorical features and clipped the number of categories for features with many unique values. We apply either an identical mapping or a logarithmic transformation for numerical features. We use the gradient boosting algorithm from PySpark ML for downstream tasks.
5.2. Downstream tasks
5.2.1. MBD: campaigning task
@@
This Subsection contains experimental results for the campaigning task in the MBD dataset in unimodal and multimodal baselines. One of the main objectives of our paper is to provide a real data benchmark to facilitate the development of multimodal algorithms. It is necessary to demonstrate that an algorithm outperforming another on our public benchmark will similarly outperform it on real data.
@@
Table 2 shows that the transaction modality is crucial in achieving accurate classification. In contrast, dialogues and geostream perform only slightly better than a random estimator when used in unimodal setting. However, as shown in Table 3, the predictive performance improved significantly in the multimodal setting by integrating additional modalities. The overall trend indicates a consistent improvement in validation metrics as more modalities are incorporated. Specifically, the multimodal late fusion approach enhances predictive accuracy by 1-1,5% when adding other data sources to the transaction stream.
Model performance comparison on private and public data.
Fig. 2 highlights a strong correlation between performance metrics on public and private datasets, with a Kendall-tau correlation coefficient of 0.94. This figure presents performance results derived from the extended experiments detailed in Appendix B (Tables 11, 13) for the campaigning task on both datasets. The private dataset is structurally equivalent to the public dataset, maintaining identical client composition, temporal coverage, and data volume, with the sole distinction being the application of anonymization. Hence, the anonymization process has minimal impact on model performance for the downstream task of campaigning. The consistency in relative ranking across both datasets (Table 3) underscores the reliability of our benchmark for advancing research in multimodal event sequence analysis.
@@
More detailed comparison of results on MBD and the private dataset and comprehensive results for each modality and all possible fusion combinations of multiple modalities are shown in Appendix B (Tables 10 - 13). Our modalities, namely, geostream, transactions, and dialogues, are denoted as Geo, Trx, and Dialogs. We use a clear notation to specify a method applied to a modality. For instance, if embeddings from the CoLES model are applied to transactions, we denote this as TrxCoLES.
5.2.2. Datafusion: Higher Education
In this subsection, we present the results of unimodal and multimodal experiments on the DataFusion dataset, as shown in Tables 2 and 3. While the dataset is small, leading to an insignificant increase in quality metrics when incorporating additional sources in multimodal settings, the results in both unimodal and multimodal configurations remain valuable as they contribute to expanding our benchmark.
5.2.3. Alphabattle: Default
In Table 2, we present the metrics for unimodal methods on the Alphabattle dataset. Interestingly, the supervised methods on both the MBD and Alphabattle datasets perform as well as, or better than, unsupervised methods. This may be attributed to the dataset size and the available labeled data.
5.2.4. Multimodal matching task
@@
In this subsection, we present the results of our proposed multimodal matching benchmark, summarized in Table 4, which includes both MBD and Datafusion datasets. Detailed results for other modalities can be found in Appendix B, in Table 8. We report Recall@1, Recall@50, and Recall@100 for each modality, measured in both directions. We also present the multimodal matching results for the DataFusion dataset. For example, in the case of the transaction and geostream pair for MBD, we compute both Trx2Geo and Geo2Trx, evaluating alignment from both perspectives. Our analysis reveals considerable variation in performance across different modality pairs. Specifically, dialogues consistently exhibit weaker matching performance than other modalities, such as transactions and geostream, demonstrating significantly stronger alignment. This disparity suggests potential limitations within the dialogue modality, indicating that it may offer less complementary or aligned information. Alternatively, the unique structure of dialogue data may necessitate more sophisticated techniques for effective integration with other modalities.
| Dataset | Modalities | Recall@1 | Recall@50 | Recall@100 |
| MBD | Trx2Geo | 0.006 0.0003 | 0.196 0.002 | 0.303 0.004 |
| Geo2Trx | 0.004 0.0003 | 0.162 0.002 | 0.262 0.004 | |
| Datafusion | Trx2Click | 0.002 0.0010 | 0.063 0.004 | 0.120 0.009 |
| Click2Trx | 0.001 0.0007 | 0.070 0.005 | 0.115 0.008 |
| Dataset | Modalities | Blending | Late Fusion | Early Fusion | |
| TabGPT | TabGPT | Supervised RNN | CrossTransformer | ||
| MBD | Trx + Geo | 0.804 0.001 | 0.800 0.001 | 0.819 0.001 | 0.815 0.001 |
| Trx + Dialog | 0.742 0.001 | 0.810 0.002 | 0.821 0.0006 | 0.821 0.002 | |
| Datafusion | Trx + Click | 0.756 0.013 | 0.766 0.011 | 0.703 0.008 | 0.735 0.010 |
5.3. Comparison of Multimodal Fusion Methods
In this subsection, we evaluate the performance of multimodal fusion techniques across multiple datasets. For this analysis, we select the best-performing models for Blending and Late Fusion to compare against Early Fusion. As shown in Table 5, Late Fusion with a Supervised RNN consistently delivers superior results on the MBD dataset, demonstrating its robustness in effectively integrating multimodal data. Early Fusion, implemented via the cross-attention from CrossTransformer (Zhang and Yan, 2023), achieves competitive performance but slightly lags behind Late Fusion, while Blending exhibits considerably weaker results. On the Datafusion dataset, Late Fusion with TabGPT is the most effective method, with Early Fusion also performing well and outperforming Late Fusion with a Supervised RNN. Blending, in contrast, consistently yields the lowest performance across all modalities. These findings highlight the effectiveness and reliability of Late Fusion for multimodal data integration while identifying Early Fusion as a promising direction for further research and optimization.
5.4. LLMs for multimodal banking data
Large Language Models, or LLMs, are now strong tools for handling various types of tabular data (Ruan et al., 2024). This includes sequential tabular data (Zhang et al., 2023). We used popular models LLAMA3.2 (Meta, 2024) and QWEN2.5 (Yang et al., 2024). These two are publicly available. We also used two in-house LLMs, Gigar4 and Gigar4 SFT (Supervised Fine-Tuning). The last two models are Russian language-specific language models. Gigar4 SFT is a version that has been fine-tuned on several math datasets. The fine-tuning of math datasets enhances their ability to process and predict numerical data accurately. We took 3B parameter versions of the models to keep inference time low while being able to process context up to 8192 tokens on a single GPU.
Experiments with LLMs are computationally expensive, so we used the MBD-small subset described in Subsection 3.1. Inference on this dataset using 3B parameters LLMs takes 12 hours on a single NVIDIA A100 GPU. Our results (Table 6) show that in transactions, relatively small language models can match the performance of specialized event-sequence models, even without extra tuning. At the same time, there is room for improvement to better account for data-specific characteristics. Geo hashes exhibit extreme diversity in values (44K unique anonymized values even for level-4 hashes), which makes it challenging for models to generalize without fine-tuning. Additionally, adding modalities increases the required context length, which is already a challenge for LLMs, given the long event sequences and the way they are encoded into text.
| Model | Trx | Trx+Geo | Geo | |
| Gigar4 | 0.7730.007 | 0.7370.006 | 0.5480.004 | |
| LLM | Gigar4 SFT | 0.7780.005 | 0.7350.014 | 0.5490.010 |
| LLAMA3.2 | 0.7730.006 | 0.7350.011 | 0.5340.016 | |
| QWEN2.5 | 0.7570.005 | 0.7270.007 | 0.5390.008 | |
| CoLES | 0.7850.006 | 0.7820.005 | 0.5680.007 | |
| TabGPT | 0.7790.007 | 0.7820.002 | 0.5940.016 | |
| Baselines | TabBERT | 0.7810.003 | 0.7890.002 | 0.6010.013 |
| Aggregation | 0.7810.009 | 0.7830.008 | 0.5300.006 | |
| Supervised RNN | 0.7450.004 | 0.7420.006 | 0.5320.013 |
6. Conclusion and Future Work
In this paper, we introduce MBD, the first large-scale, publicly available multimodal banking dataset (Table 1), comprising sequential data from over 2 million clients, including transactions, geo-locations, and support dialogues. MBD ensures compliance with data protection regulations, enabling research on financial data while maintaining privacy. Our findings show that anonymization has minimal impact on algorithm performance (Table 3, Fig. 2), making the dataset suitable for real-world deployment. Additionally, excluding sensitive attributes (e.g., gender, age, race) mitigates bias, fostering ethical AI development.
MBD, along with Datafusion and Alphabattle, establishes a benchmark for key downstream tasks, facilitating the development of scalable multimodal and unimodal algorithms for event sequence prediction. We implemented the state-of-the-art models in banking domains, demonstrating that even basic multimodal fusion outperforms single-modal baselines (Table 3). Further, Late and Early Fusion experiments (Table 5) highlight the potential for improving model performance by effectively capturing cross-modal interactions. We also assess LLMs in unimodal and multimodal settings, comparing them with domain-specific approaches (Table 6). Given MBD’s scale and practical relevance, even modest performance gains can yield significant financial benefits.
Future iterations will incorporate additional data sources (e.g., clickstreams, enriched dialogue features) and expand downstream tasks to financial product recommendations, customer churn prediction, fraud detection, and credit risk assessment. Extending the dataset’s time range will further enable longitudinal studies and trend analysis
Appendix A Dataset Statistics and data samples
In this section, we present data samples from various modalities and a histogram of sequence lengths. Table 7 provides a list of attributes for the transactional modalities, while Figure 3 illustrates the distribution of sequence lengths for these modalities. Additionally, Figure 6 shows a sample of geographical data, and Figure 7 displays the distribution of sequence lengths for geostream modalities. For dialogue data, a sample is presented in Figure 4, with the distribution of sequence lengths for dialogue events shown in Figure 5. Figure 8 demonstrates a sample of data for product purchases. Here, we observe that geographical and transactional modalities have long tails in the histogram of distributions.